# CayleyPy RL: Pathfinding and Reinforcement Learning on Cayley Graphs

### v1

@@ Abstract.
This paper is the second in a series of studies on developing efficient artificial intelligence-based approaches to pathfinding on extremely large graphs (e.g. $10^{70}$ nodes) with a focus on Cayley graphs and mathematical applications. The open-source CayleyPy project is a central component of our research. The present paper proposes a novel combination of a reinforcement learning approach with a more direct diffusion distance approach from the first paper. Our analysis includes benchmarking various choices for the key building blocks of the approach: architectures of the neural network, generators for the random walks and beam search pathfinding. We compared these methods against the classical computer algebra system GAP, demonstrating that they ”overcome the GAP” for the considered examples. As a particular mathematical application we examine the Cayley graph of the symmetric group with cyclic shift and transposition generators. We provide strong support for the OEIS-A186783 conjecture that the diameter is equal to n(n-1)/2 by machine learning and mathematical methods. We identify the conjectured longest element and generate its decomposition of the desired length. We prove a diameter lower bound of n(n-1)/2-n/2 and an upper bound of n(n-1)/2+ 3n by presenting the algorithm with given complexity. We also present several conjectures motivated by numerical experiments, including observations on the central limit phenomenon (with growth approximated by a Gumbel distribution), the uniform distribution for the spectrum of the graph, and a numerical study of sorting networks. To stimulate crowdsourcing activity, we create challenges on the Kaggle platform and invite contributions to improve and benchmark approaches on Cayley graph pathfinding and other tasks.
@@ 1.1. Broader context
Deep learning has revolutionized various fields of research and was recognized by several Nobel Prizes in 2024. In recent years, machine learning has been emerging as ”a tool in theoretical science” [Douglas22], leading to several noteworthy applications to mathematical problems: [Charton19, Davies21, Bao23, Romera24, Coates23, Hayat25, Williamson24, Gukov24, Kohli25].
The paper presents advancements in applying artificial intelligence methods to the mathematical problem of Cayley graph pathfinding, which is equivalent to decomposing a group element into a product of generators. It is the second work in a series devoted to the CayleyPy project to develop efficient machine learning-based approaches and an open source library to deal with extra large graphs (e.g. $10^{70}$) with a focus on finite Cayley graphs and mathematical applications.
From a broader perspective, pathfinding is a specific case of the planning problem, where one must determine a sequence of actions to transition between two given states. Similar challenges arise in robotics and games like chess or Go, where the objective is to plan moves that lead from the initial position to a winning position. Mathematically, such problems are modeled as pathfinding on graphs (state transition graphs), where nodes represent possible states, and edges correspond to state transitions based on actions or moves. The planning task then reduces to finding a path from a given initial node to one or more target nodes. The breakthrough works AlphaGo and AlphaZero [AlphaGo16], [AlphaZero17] by Google DeepMind have demonstrated that machine learning can significantly outperform previously known methods. They serve as both a prototype and an inspiration for many developments, including ours. In a nutshell, the method consists of two main components: a neural network trained to suggest moves from any given state, and a graph pathfinding algorithm that helps to find a path even when the neural network’s predictions are not sufficiently precise. The core idea is similar across most approaches, but achieving optimal performance for each task requires analyzing and benchmarking various neural networks and graph pathfinding methods. One of the goals of this work is to present such an analysis for some particular Cayley graphs.
1.2. Present paper: approach, benchmarks, mathematical theorems and conjectures
@@
The paper presents both AI contributions and mathematical results and conjectures. Here, we work with a particular Cayley graph defined by the L - left cyclic shift, R - right cyclic shift and X - eXchange of positions $0,1$, it will be denoted LRX following OEIS-A186783. Its image for $S_{5}$ is shown on figure 1.
@@
Developments of AI methodology. We propose a modified deep Q-learning (DQN) method for the case of the graph pathfinding task which resolves the so-called sparse reward problem of the standard DQN approach for graph pathfinding. The proposal combines DQN with the more simple diffusion distance approach from our previous paper  [Chervov25]. An analysis and benchmarks of these approaches are presented, including tests of various neural network architectures such as transformers, convolutional networks, and perceptrons. It is demonstrated that these methods allow finding paths for $S_{n}$ up to $n$ around 30 - for the particular class of generators (LRX) which are in the main focus of the present paper. We demonstrate that they ”overcome the GAP” - outperform classical computer algebra system GAP on the same task which can work only up to $n$ around 20.
@@
Efficient pytorch implementation. We pay significant attention to develop an original highly optimized code, the current version is specific to Cayley graphs of permutation groups (matrix groups will be supported later). We propose many technical solutions for fast parallel processing of large amounts of permutations, including product computations, efficient hashing, and extracting unique states. The code works on both CPU and GPU without modification.
@@
One line of code bringing orders of magnitudes improvement. We report a curious finding - a single-line modification in the code of our main graph pathfinding module (beam search), dramatically improves performance. This enhancement increases the feasible pathfinding size from symmetric groups $S_{n}$ with $n$ to $n$ around 100 (and apparently even further with more hardware). Unfortunately, this improvement is currently specific to the particular LRX Cayley graphs considered in this paper, but we hope it can be generalized.
Mathematical contributions.
LRX Cayley graphs. Diameter conjecture OEIS-A186783. In the present paper, we focus on a specific family of Cayley graphs for the symmetric groups $S_{n}$ those generated by the two cyclic shifts (L) - left, (R) - right and the transposition X=$(0,1)$ of the first two positions. Surprisingly, according to OEIS-A186783, it is an open conjecture that the diameter of the Cayley graph is $n(n-1)/2$ for $S_{n}$. These generators and Cayley graphs will be denoted LRX as in OEIS.
@@
The longest element. By implementing an efficient brute-force breadth-first search algorithm, we traverse the full LRX graphs for $n\leq 14$, confirming that the longest element has a length of $n(n-1)/2$, as predicted. Moreover, we observe that this element is always unique and follows a clear pattern - it is a product of transpositions: $(0,1)(n-1,2)(n-2,3)(n-3,4)...$, so we expect this for all $n$.
@@
AI and crowd-sourcing to support the conjecture. For that particular element, we have launched various versions of our pipelines and have consistently found that the decomposition is never shorter than $n(n-1)/2$. We organized a challenge on the Kaggle platform giving that particular element for $n=2...200$ and requesting decompositions of the shortest possible length. Once again, no participant found a decomposition shorter than $n(n-1)/2$, thereby confirming the expectation.
@@
Rigorous lower and upper bounds. We prove lower bound $n(n-1)/2-n/2-1$ by a combinatorial argument (improving $n(n-1)/3$ [Babai89]); afterwards, we prove the diameter upper bound $n(n-1)/2+3n$ by the algorithm with such complexity (improving $3/2n^{2}$  [Kuppili20]). We also develop another algorithm which empirically shows better complexity: $n(n-1)/2+n/2$, but lacks the rigorous proof for such an estimate. Finally, we present explicit decomposition for the conjecturally longest element of the desired length $n(n-1)/2$.
@@
Central limit phenomena for growth - Gumbel distribution. Computations suggest the following conjecture: for large $n$, the growth function of LRX graphs follows an asymmetric Gumbel distribution. This is in the vein of field-shaping works by P. Diaconis et. al. [Diaconis77, Diaconis88, Diaconis17], who demonstrated that the growth of the neighbor transposition graph and many related statistics asymptotically follow the Gaussian normal distribution.
@@
We study numerically the spectrum of LRX graph and observe surprisingly uniform distribution of eigenvalues, we also analyze statistics of the paths from the longest element to the identity (”sorting networks”) we observe numerically that pattern differs from the famous ”sine curves” for Cayley graph of neighbor transpositions [Virág]; numerical analysis of random walks shows consistency with results of Diaconis [Diaconis93] on mixing time for LRX Cayley graph.
We also present similar analysis for the Schreier coset graph for the action of the LRX generators on binary strings consisting of $n/2$ zeros and ones (for even $n$).
@@
In summary, our analysis demonstrates that computational experiments, especially those involving AI, can be very helpful in advancing mathematical research on Cayley graphs. AI methods significantly outperform classical methods of computer algebra systems such as GAP, performing well for $S_{n}$ with $n$ around 30, and with a small addition of prior knowledge up to 90, while classical methods perform well only up to 20.
@@
1.3. Finite Cayley graphs: pathfinding techniques, results and open problems
@@
Let us briefly summarize existing techniques for Cayley graphs pathfinding, and remind more general context on Cayley graphs research.
@@
Optimal pathfinding is typically NP-hard. Finding the shortest paths on generic finite Cayley graphs is an NP-hard problem [Goldreich81] (even P-space complete [Jerrum85]). As is the case for many specific group families, such as NxNxN Rubik’s Cube groups [Demaine17] and others [Bulteau15]. For particular groups it is hard to develop practical optimal algorithm e.g. standard 3x3x3 Rubik’s cube the first optimal algorithm has been proposed only 1997 [Korf97]. It required 192 Gigabytes of precomputed data and 12-24 hours to solve single cube (methods were later improved  [^1] M.Reid: https://www.cflmath.com/Rubik/optimal_solver.html, H.Kociemba 2021: https://pypi.org/project/RubikOptimal/, Andrew Skalski: https://github.com/Voltara/vcubem , [Rokicki14] to several cubes per second). To develop practical optimal solver for 4x4x4 Rubik’s cube - is a challenge.
@@
Non-optimal pathfinding: Schreier-Sims algorithm, GAP implementation, non-effectivity for large groups. The Schreier-Sims algorithm [Sims70],[Knuth91] (or its randomized versions) is the method typically used for decomposing group element into product of generators for permutation groups. It is implemented in GAP computer algebra system. However its outputs ”are usually exponentially long” [Shamir89] and that forbids practical computations for large groups (typically the sizes beyond $10^{30}-10^{40}$ are out of reach - examples and analysis is given below).
@@
Non-optimal pathfinding: ”small support” Kaggle Santa 2023 methods (generality issue). The methods which can deal with extremely large groups (e.g. size $>10^{1000}$ like 33x33x33 Rubik’s cube) were demonstrated by top participants of the recent challenge organized on the Kaggle platform - Santa 2023. However the generality of that approach is unclear.That is contrast to AI methodology which is from the beginning is general not even restricted to Cayley graphs. The idea of ”Santa methods” is quite noteworthy: one should first find ”small support elements” achievable from original generators and take them as new generators, then one simply runs beam search with Hamming distance heuristics to find path. Method becomes effective since for small support generators Hamming distance and true graph distance are quite related by obvious reason. ”Small support” are those permutations which change only small number of positions. They are well-known for Rubik’s cube solvers - ”commutator” or ”Pif-Paf” moves [Mulholland16], M.Brandenburg post. In mathematics small support elements play an important role in deep works by Babai, Seress, Helfgott, et.al. [Babai88, Babai04, Seress14, Helfgott14, Helfgott19] (slides: A.Seress,H.Helfgott ), however it is unclear whether one can effectively find small support elements from given generators, thus making unclear the generality of ”small support methods”.
@@
Non-optimal pathfinding: algorithms for specific families of generators. For some particular choices of generators there are algorithmic solvers which can effectively find non-optimal solution. It seems unlikely that one can produce similar algorithms for any generators. The prototypical example is bubble sort algorithm, which solves the problem for the neighbor transpositions generators of $S_{n}$. Algorithmic solvers exist for Rubik’s cube of arbitrary sizes and for other puzzles related groups. The other families of generators and algorithms are important in bioinformatics for estimation of the evolutionary distance. These generators are related to flips of subsequences - see e.g. [Pevzner95, Pevzner99, Bulteau19]. Related example is pancake sorting (or prefix sorting) problem with the first algorithm proposed by Bill Gates et.al. [Gates79].
@@
AI-methods for pathfinding. To the best of our knowledge there is no systematic effort for Cayley graph pathfinding with AI methods except our CayleyPy project. However for some particular cases for example Rubik’s cube there exists a number of approaches, e.g. based on genetic algorithms: [Świta23].The most notable are: DeepCube [McAleer19, Agostinelli19, Agostinelli24, Agostinelli24q] - the first AI based successful solver for 3x3x3 cube (website), the second one: EfficientCube [Takano21], some other: [TakanBrunetto17, Johnson21, Amrutha22, Noever21, Chasmai21, Bedaywi23, Pan21] which not achieve complete solution. Noteworthy idea [Pan21] is to combine neural networks with representations theory of the symmetric group — neural net predicts the coefficients of the non-abelian Fourier transform for the distance function. The rationale is: observed sparsity (bandlimitedness) of Fourier transform of the common distance functions on $S_{n}$  [Swan17]. The recent seminal achievement from S.Gukov’s team:  [Gukov24] creates effective AI based pathfiding approach for infinite group of Andrews-Curtis moves and resolves the Akbulut-Kirby conjecture remained opened for 39 years. AI-methods are also used for pathfinding in the context of planning movements in obstacle-rich environments and road networks it is related field, but with different focus [Leskovec22, Yakovlev23]
@@
Cayley graphs: applications, diameters, random walks, open conjectures. Let us give a more general context on Cayley graph research. Cayley graphs are fundamental in group theory [Gromov93],[Tao15], and have various applications: bioinformatics  [Pevzner95, Pevzner99, Bulteau19]; processors interconnection networks [Akers89, Cooperman91, Heydemann97];in coding theory coding theory and cryptography [Wigderson06, Zémor94, Petit11]; in quantum computing [Kohli24, Sarkar24, Vidick23, Acevedo06, Gromada22], etc.
@@
Two focuses of pure mathematical research on finite Cayley graphs are related to diameters estimation and the behavior of random walks. In the case of the symmetric group $S_{n}$ there are two open conjectures which are easy to formulate, but somewhat representative for the field.
Babai-like conjecture: diameter of $S_{n}$ is not more than $O(n^{2})$ for any choices of generators (see e.g. H.Helfgott’s surveys:  [Helfgott14],[Helfgott19],[Helfgott15] ;
@@
Diaconis  [Diaconis13] conjecture: the mixing time for random walks is not more than $O(n^{3}log(n))$ (again for any choices of generators).
@@
The first one can be thought as a particular case of the Babai conjecture [Babai92], [Tao15] which predicts that diameter of simple groups is not as large as one may expect: $O(log^{c}|G|)$ for some absolute $c$ — in contrast to abelian groups where it can be $O(|G|)$. Thus conjecture represents an intriguing connection between algebraic properties of a group (which does not depend on the choice of generators) - being simple and the property of Cayley graphs - having short diameter (which is generator dependent property).

### v2

@@ Abstract.
This paper is the second in a series of studies on developing efficient artificial intelligence-based approaches to pathfinding on extremely large graphs (e.g., $10^{70}$ nodes) focusing on Cayley graphs and mathematical applications. The open-source CayleyPy project is a central component of our research. The present paper proposes a novel combination of a reinforcement learning approach with a more direct diffusion distance approach from the first paper. Our analysis includes benchmarking various choices for the key building blocks of the approach: neural network architectures, random walk generators, and beam search pathfinding. We compared these methods against the classical computer algebra system GAP, demonstrating that they ”overcome the GAP” for the considered examples. We examine the Cayley graph of the symmetric group with cyclic shift and transposition generators as a particular mathematical application. We strongly support the OEIS-A186783 conjecture that the diameter equals n(n-1)/2 by machine learning and mathematical methods. We identify the conjectured longest element and generate its decomposition of the desired length. We prove a diameter lower bound of n(n1)/2n/21n(n-1)/2-n/2-1 and an upper bound of n(n1)/2+3nn(n-1)/2+3n by presenting the algorithm with given complexity. We also present several conjectures motivated by numerical experiments, including observations on the central limit phenomenon (with growth approximated by the Gumbel distribution), the uniform distribution for the spectrum of the graph, and a numerical study of sorting networks. We create challenges on the Kaggle platform to stimulate crowdsourcing activity and invite contributions to improve and benchmark approaches to Cayley graph pathfinding and other tasks.
@@ 1.1. Broader context
Deep learning has revolutionized various fields of research and was recognized with several Nobel Prizes in 2024. In recent years, machine learning has been emerging as ”a tool in theoretical science” [Douglas22], leading to several noteworthy applications to mathematical problems: [Charton19, Davies21, Bao23, Romera24, Coates23, Hayat25, Williamson24, Gukov24, Kohli25].
The paper presents advancements in applying artificial intelligence methods to the mathematical problem of Cayley graph pathfinding, equivalent to decomposing a group element into a product of generators. It is the second work in a series devoted to the CayleyPy project to develop efficient machine learning-based approaches and an open-source library to deal with extra-large graphs. (e.g. $10^{70}$) with a focus on finite Cayley graphs and mathematical applications.
From a broader perspective, pathfinding is a specific case of the planning problem, where one must determine a sequence of actions to transition between two given states. Similar challenges arise in robotics and games like chess or Go, where the objective is to plan moves that lead from the initial position to a winning position. Mathematically, such problems are modeled as pathfinding on graphs (state transition graphs), where nodes represent possible states, and edges correspond to state transitions based on actions or moves. The planning task then reduces to finding a path from a given initial node to one or more target nodes. The breakthrough works AlphaGo and AlphaZero [AlphaGo16], [AlphaZero17] by Google DeepMind have demonstrated that machine learning can significantly outperform previously known methods. They serve as both a prototype and an inspiration for many developments, including ours. The method consists of two main components: a neural network trained to suggest moves from any given state and a graph pathfinding algorithm that helps to find a path even when the neural network’s predictions are not sufficiently precise. The core idea is similar across most approaches, but achieving optimal performance for each task requires analyzing and benchmarking various neural networks and graph pathfinding methods.
Graph Group RL
vertex / node permutation / element state
edge generator action
path word trajectory
shortest path shortest word maximal return
diameter longest element ?
Table 1. Comparison of graph theory, algebraic and reinforcement learning terminology.
One of the goals of this work is to apply these methods to a particular class of Cayley graphs. As in our previous paper [Chervov25], we choose a set of generators $a_{1}^{\pm 1},\ldots,a_{k}^{\pm 1}\in S_{n}$. All finite products (words) of these elements form a subgroup $G\subset S_{n}$ which we identify with its Cayley graph. Due to vertex-transitivity of Cayley graphs, the pathfinding problem can be reduced to searching a path from some element $g\in G$ to the identity permutation $e$; this is the same as finding a decomposition of $g$ into a word $\prod\limits_{j=1}^{m}a_{i_{j}}^{\pm 1}$, $1\leq i_{j}\leq k$. The problem of finding the shortest path (or word) can also be reformulated in terms of a generic reinforcement learning approach discussed in Section 2.1. A set of dictionary between RL, graphs, and group is presented in Table 1.
1.2. Present case study - LRX Cayley graph.
@@
The paper presents both AI contributions and mathematical results and conjectures. We use LRX Cayley graphs as primary benchmark in the subsequent discussion.
LRX Cayley graphs. For each $n\in\mathbb{N}$ the symmetric group $S_{n}$ can be generated by three permutations:
  1. (1)
    left cyclic shift $L=(1\ 2\ 3\ \ldots\ n-1\ 0)$;
  2. (2)
    right cyclic shift $R=L^{-1}=(n-1\ 0\ 1\ 2\ \ldots\ n-3\ n-2)$;
  3. (3)
    transposition (eXchange) of the first two positions $X=X^{-1}=(0,1)$.
These generators and the corresponding Cayley graph will be denoted LRX following OEIS-A186783. Visualization (Fig. 1) of the LRX Cayley graph for $S_{5}$ reveals its complicated structure. LRX Cayley graph can be considered as a relative of the Oval Track (or TopSpin) and Hungarian Rings puzzles (Fig. 2).
Figure 2. LRX and related mechanical puzzles
Conjectures on diameters of LRX and LX (V.M. Glushkov’s 1968 problem). It is conjectured OEIS-A186783 that diameters of LRX are n(n1)/2n(n-1)/2, that conjecture will be in the focus of the present paper. The case of two generators just and L and X (not inverse closed) goes back at least to 1968 influential paper by “one of the founding fathers of Soviet cybernetics” V. M. Glushkov [Glushkov68], which has been much studied (survey: [GlukhovZubov99] pages 18-21), but not resolved. The conjecture on the diameter in that case is presented in our subsequent paper: [CayleyPyGrowth].
Not exactly in the present scope, but worth mentioning: there was a long-standing problem resolved positively recently [SawadaWilliams19] of whether the directed graph generated by L and X (also denoted Sigma and Tau) has a hamiltonian path. D.Knuth assigned to that problem complexity 48 out of 50, it was first articulated in [NijenhuisWilf75].
1.3. AI contributions.
@@
Development of AI methodology. We suggest a modified deep Q-learning (DQN) method for the case of the graph pathfinding task, which resolves the so-called sparse reward problem of the standard DQN approach for graph pathfinding. The proposal combines the DQN and the diffusion distance approach from [Chervov25] and presents an analysis and benchmarks of these techniques, including tests of various neural network architectures such as MLP, CNN, and transformers. These methods have been shown to find paths for LRX Cayley graphs up to $n$ around 40 and with small modification (X-trick, which is specfic to LRX) even beyond 100. We demonstrate that they ”overcome the GAP”: outperform the classical computer algebra system GAP on the same task, which can work only up to $n$ around 20, also producing much more optimal paths and working faster.
@@
Efficient PyTorch implementation. We devote significant attention to developing an original, highly optimized code. The current version is specific to Cayley graphs of permutation groups (matrix groups will be supported later). We propose many technical solutions for fast parallel processing of large amounts of permutations, including product computations, efficient hashing, and extracting unique states. The code works on both CPU and GPU without modification.
@@
X-trick. We report a curious finding: a single-line modification (which we call X-trick) in the code of our main graph pathfinding module (beam search) dramatically improves performance. This enhancement increases the feasible pathfinding size to $n$ around 100 (apparently even further with more powerful hardware). Unfortunately, this improvement is currently specific to the particular LRX Cayley graphs considered in this paper, but we hope it can be generalized. Moreover, we provide a mathematical counterpart of that finding - X-trick modified random walks are mixing much faster - see discussion below.
1.4. Mathematical contributions.
We present several conjectures about LRX Cayley graphs which could leverage a more general understanding of the properties of that Cayley graph and its relatives.
@@
Diameter conjecture and the longest element. Surprisingly, according to OEIS-A186783, it is an open conjecture that the diameter of the Cayley graph is n(n1)/2n(n-1)/2 for $S_{n}$. By implementing an efficient brute-force breadth-first search algorithm, we traverse the full LRX graphs for $n\leq 15$, confirming that the longest element has a length of n(n1)/2n(n-1)/2, as predicted. Moreover, we observe that this element $\ell_{n}$ is always unique and follows a clear pattern
$\ell_{n}=(1\ 0\ n-1\ n-2\ \ldots\ 3\ 2)=(0,1)(n-1,2)(n-2,3)\ldots=R^{2}e^{\leftarrow},$ (1)
where $e^{\leftarrow}=(n-1\ n-2\ \ldots\ 2\ 1\ 0)$ is reversed identity permutation. We expect (1) holds for all $n$. Interestingly, the longest element $\ell_{n}$ is equal not to the reversed identity permutation itself but to the element obtained by two right shifts from it.
@@
AI and crowd-sourcing to support the conjecture. For the longest element $\ell_{n}$, we have launched various versions of our pipelines, and have consistently found that a decomposition is never shorter than n(n1)/2n(n-1)/2. We organized a challenge on the Kaggle platform suggesting to decompose $\ell_{n}$, $n=2,\ldots,200$, into the shortest possible word. Once again, no participant found a decomposition shorter than n(n1)/2n(n-1)/2, thereby confirming the expectation.
@@
Rigorous lower and upper bounds for the diameter. We prove the lower bound n(n1)/2n/21n(n-1)/2-n/2-1 by a combinatorial argument (improving n(n1)/3n(n-1)/3 [Babai89]); later, we prove the upper bound n(n1)/2+3nn(n-1)/2+3n by the algorithm with such complexity (improving $\frac{3}{2}n^{2}$ [Kuppili20]). We also develop another algorithm that empirically shows better complexity n(n1)/2+n/2n(n-1)/2+n/2, but it lacks a rigorous proof for such an estimate. Finally, we present an explicit decomposition for the conjecturally longest element $\ell_{n}$ of the desired length n(n1)/2n(n-1)/2.
@@
Central limit phenomena for growth. Computations suggest the following conjecture: for large $n$, the growth function of LRX graphs follows a left-skewed Gumbel distribution. This statement is in the vein of field-shaping works by P. Diaconis et. al. [Diaconis77, Diaconis88, Diaconis17], who demonstrated that the growth of the neighbor transposition graph and many related statistics asymptotically follows the Gaussian normal distribution. However we observe that skewness on the growth distribution for LRX does not tend to zero, thus Gaussian is not a choice, while Gumbel provides a good numerical fit. We also present a similar analysis for the Schreier coset graph for the action of the LRX generators on binary strings consisting of $n/2$ zeros and ones (for significantly larger range of $n$).
Random walks and mixing time. P.Diaconis and L.Saloff-Coste [Diaconis93] obtained an important result that mixing time for LRX is between $O(n^{3})$ and $O(n^{3}log(n))$. We support their result numerically and complement it by considering two other types of random walks: non-backtracking and X-trick modified. Both show faster mixing, however non-backtracking is only faster by a constant factor, while X-trick leads to around $O(n^{2})$ mixing. Which particially explains its effectivity in beam search.
@@
Spectrum and sorting networks. We study numerically the spectrum of the LRX graph and observe a surprisingly uniform distribution of eigenvalues. Investigating statistics of the paths from $\ell_{n}$ to $e$ (”sorting networks”), we observe numerically that the pattern differs from the famous ”sine curves” for Cayley graph of neighbor transpositions [Virág].
@@
In summary, our analysis demonstrates that computational experiments, especially those involving AI, can be very helpful in advancing mathematical research on Cayley graphs. AI methods significantly outperform classical methods of computer algebra systems such as GAP, performing well for $S_{n}$ with $n$ around 40 and with a small addition of prior knowledge up to 90, while classical methods perform well only up to 20.
@@
1.5. Review. Finite Cayley graphs: pathfinding techniques, results and open problems
@@
Let us briefly summarize existing techniques for Cayley graphs pathfinding, and provide more general context on Cayley graphs research.
@@
Optimal pathfinding is typically NP-hard. Finding the shortest paths on generic finite Cayley graphs is an NP-hard problem [Goldreich81] (even P-space complete [Jerrum85]). As is the case for many specific group families, such as $N\times N\times N$ Rubik’s Cube groups [Demaine17] and others [Bulteau15]. It is hard to develop a practical optimal algorithm for particular groups, for instance, the first optimal algorithm for the standard $3\times 3\times 3$ Rubik’s cube has been proposed only in 1997 [Korf97]. It required 82 Megabytes of precomputed data and 12-24 hours to solve a single cube (later improved  [^1] M.Reid: https://www.cflmath.com/Rubik/optimal_solver.html, H.Kociemba 2021: https://pypi.org/project/RubikOptimal/, Andrew Skalski: https://github.com/Voltara/vcubem , [Rokicki14] to several cubes per second). Developing a practical optimal solvers for a $4\times 4\times 4$ Rubik’s cube is challenging open problem.
@@
Non-optimal pathfinding: Schreier-Sims algorithm, GAP implementation, non-effectivity for large groups. The Schreier-Sims algorithm [Sims70] (and its randomized version [Knuth91]) is the method typically used for decomposing group elements into products of generators for permutation groups. It is implemented in the GAP computer algebra system. However, its outputs ”are usually exponentially long” [Shamir89], and that forbids practical computations for large groups (typically the sizes beyond $10^{30}$$10^{40}$ are out of reach – examples and analysis are given below).
@@
Non-optimal pathfinding: ”small support” Kaggle Santa 2023 methods (generality issue). The methods which can deal with extremely large groups (e.g., $10^{1000}$ in size like $33\times 33\times 33$ Rubik’s cube) were demonstrated by top participants of the recent Kaggle challenge Santa 2023. However, the generality of that approach is unclear. That is in contrast to AI methodology, which is, from the beginning general not even restricted to Cayley graphs. The idea of ”Santa methods” is quite noteworthy: one should first find ”small support elements” achievable from original generators and take them as new generators, then one simply runs beam search with Hamming distance heuristics to find path. This method becomes effective since for small support generators, Hamming distance, and true graph distance is quite related for obvious reasons. ”Small support” is those permutations that change only a small number of positions. They are well known for Rubik’s cube solvers – ”commutator” or ”Pif-Paf” moves [Mulholland16], M.Brandenburg post. In mathematics, small support elements play an important role in deep works by Babai, Seress, Helfgott, et.al. [Babai88, Babai04, Seress14, Helfgott14, Helfgott19] (slides: A.Seress,H.Helfgott). However, the question whether one can effectively find small support elements from given generators remains open, thus making the generality of ”small support methods” unclear.
@@
Non-optimal pathfinding: algorithms for specific families of generators. Some algorithmic solvers can effectively find non-optimal solutions for some particular generator choices. It seems unlikely that one can produce similar algorithms for any generators. The prototypical example is the bubble sort algorithm, which solves the problem for the neighbor transposition generators of $S_{n}$. Algorithmic solvers exist for Rubik’s cube of arbitrary sizes and other puzzle-related groups. The other families of generators and algorithms are important in bioinformatics for estimating the evolutionary distance. These generators are related to flips of subsequences (see, e.g., [Pevzner95, Pevzner99, Bulteau19]). A related example is pancake sorting (or prefix sorting) problem with the first algorithm proposed by Bill Gates et.al. [Gates79].
@@
AI-methods for pathfinding. To the best of our knowledge, there is no systematic effort for Cayley graph pathfinding with AI methods except for our CayleyPy project. However, for some particular cases, for example, Rubik’s cube, there are several approaches, e.g., based on genetic algorithms: [Świta23].The most notable are: DeepCube [McAleer19, Agostinelli19, Agostinelli24, Agostinelli24q] — the first AI based successful solver for $3\times 3\times 3$ cube (website), the second one: EfficientCube [Takano21], some other: [TakanBrunetto17, Johnson21, Amrutha22, Noever21, Chasmai21, Bedaywi23, Pan21] which not achieve complete solution. Noteworthy idea [Pan21] is to combine neural networks with the representation theory of the symmetric group — neural net predicts the coefficients of the non-abelian Fourier transform for the distance function. The rationale is: observed sparsity (bandlimitedness) of Fourier transform of the common distance functions on $S_{n}$  [Swan17]. The recent seminal achievement from S.Gukov’s team:  [Gukov24] creates an effective AI-based pathfinding approach for an infinite group of Andrews-Curtis moves and resolves the Akbulut-Kirby conjecture remained open for 39 years. AI methods are also used for pathfinding in the context of planning movements in obstacle-rich environments and road networks  [Leskovec22, Yakovlev23].
After the first version of the present work, several new contributions appeared: [Douglas25] proposed a novel diffussion-like approach to that problem, considered several classes of permutation and matrix groups and also demonstrated that the deep learning approach works successfully not only on Rubik’s cube, but on general Cayley graphs up to certain sizes. [Ziarko25] provided solution of 333 Rubik’s cube based on approach of latent representations.
@@
Cayley graphs: applications, diameters, random walks, open conjectures. Let us provide more general context on Cayley graph research. Cayley graphs are fundamental in group theory [Gromov93],[Tao15], and have various applications: bioinformatics  [Pevzner95, Pevzner99, Bulteau19]; processors interconnection networks [Akers89, Cooperman91, Heydemann97]; in coding theory and cryptography [Wigderson06, Zémor94, Petit11]; in quantum computing [Kohli24, Sarkar24, Vidick23, Acevedo06, Gromada22], etc.
@@
Pure mathematical research on finite Cayley graphs focuses on diameter estimation and the behavior of random walks. In the case of the symmetric group $S_{n}$, two open conjectures are easy to formulate, but somewhat representative of the field:
  • Babai-like conjecture: for any choices of generators diameter of $S_{n}$ is $O(n^{2})$ (see, e.g.,  [Helfgott14],[Helfgott19],[Helfgott15]);
@@
  • Diaconis  [Diaconis13] conjecture: the mixing time for random walks is $O(n^{3}\log n)$ (again for any choices of generators).
@@
The first one can be thought of as a particular case of the Babai conjecture [Babai92], [Tao15], which predicts that diameter of simple groups is not as large as one may expect: $O(\log^{c}|G|)$ for some absolute $c$ — in contrast to abelian groups where it can be $O(|G|)$. This conjecture represents an intriguing connection between an algebraic property of a group — simplicity, which is independent of the choice of generators — and a property of its Cayley graphs — having small diameter, which does depend on the generators.